# Proposal for Improving Google A2A Protocol: Safeguarding Sensitive Data in Multi-Agent Systems

### v2

@@ 6 Prompt Injection in LLM-Based Agents
Figure 4: Prompt Injection Proccess Via N8N Platform - Agent 2
@@
The findings from our experiment, along with the insights from Palo Alto article, reveal a dynamic interplay between advancing defensive measures and evolving attack strategies. As techniques to safeguard AI systems against prompt injection improve, such as enhanced input filtering or behavioral monitoring, attackers adapt by devising increasingly sophisticated methods to bypass these protections. This ongoing escalation suggests that relying solely on technical defenses may prove inadequate in the long term. Instead, the most robust strategy for safeguarding sensitive information emerges as a preventative one: refraining from sharing such data with AI agents entirely, as we propose at this paper. By excluding confidential information from AI systems, the risk of its exposure through prompt injection is effectively eliminated, aligning with the cybersecurity principle of minimizing data exposure. This conclusion not only reflects the practical lessons derived from the experiment but also emphasizes the necessity of adopting proactive and stringent security practices in the design and deployment of AI technologies. As AI continues to permeate critical sectors, ensuring its resilience against prompt injection attacks will require a concerted effort to balance functionality with security, prioritizing the protection of sensitive data above all.
7 Implementation Considerations
Implementing the proposal requires technical and organizational steps:
  • Token Management: Define processes for issuing, verifying, and revoking short-lived tokens, as recommended in [25].
  • Authentication Flows: Develop SCA flows, including two-factor or biometric options, per PSD2 requirements.
  • Protocol Extensions: Add message types or task states for sensitive data handling and consent requirements.
  • Documentation and Guidelines: Provide detailed documentation on sensitive data handling, per [42].
  • Compatibility Testing: Test enhancements with partners like PayPal and Cohere to ensure interoperability, as suggested in [42].
8 Explicit User Consent Orchestration via
USER_CONSENT_REQUIRED
Adding a new enumeration member 111 A full code sample is available at: https://github.com/yedidel/A2A :
    USER_CONSENT_REQUIRED = "user-consent-required"
The integration of USER_CONSENT_REQUIRED into the TaskState enum elevates the protocol’s control over task execution. Unlike existing states such as INPUT_REQUIRED, which merely indicate that additional information is needed, this new state explicitly denotes that a task is paused until the end-user provides affirmative consent. Embedding this state into the task lifecycle ensures that AI agents and back-end services stop sensitive operations like booking travel, executing financial transactions, or sharing personal data until a human user explicitly approves. As depicted in the diagram shown in Figure 5
Figure 5: Vacation Booking Flow, Integrated With The USER_CONSENT_REQUIRED State
This explicit state enables precise front-end handling: upon encountering consent required, the UI can display a dedicated consent prompt, ensuring user awareness and control. Once consent is granted, the task transitions to IN_PROGRESS or an equivalent active state. This mechanism not only enhances user agency and trust but also bolsters compliance with stringent regulatory frameworks (e.g., GDPR, PSD2), by providing verifiable evidence that the user approved the action before execution.
9 Conclusion
Google’s A2A protocol offers a robust foundation for secure agent communication but requires enhancements to handle sensitive data effectively. As evidenced in “AI Agents Under Threat” [20], “AI Agents Meet Blockchain” [32], “Multi-Agent Security Tax” [42], and “AgNet” [25], security and privacy challenges in multi-agent systems demand solutions balancing robust protection with efficiency. This proposal introduces seven enhancements: short-lived tokens, SCA, granular scopes, explicit consent, direct data transfer, multitransaction approval, and payment standard compliance to improve security, privacy, and trust. The vacation booking example illustrates how these enhancements reduce risks and enhance user experience. We recommend Google implement this proposal to strengthen A2A’s position as a leading standard for agent communication.

### v3

@@ VI Case Study: Prompt Injection Attack
The findings from our experiment, along with the insights from [65], reveal a dynamic interplay between advancing defensive measures and evolving attack strategies. As techniques to safeguard AI systems against prompt injection improve, such as enhanced input filtering or behavioral monitoring, attackers adapt by devising increasingly sophisticated methods to bypass these protections. This ongoing escalation suggests that relying solely on technical defenses may prove inadequate in the long term. Instead, the most robust strategy for safeguarding sensitive information emerges as a preventative one: refraining from sharing such data with AI agents entirely, as we propose at this paper. By excluding confidential information from AI systems, the risk of its exposure through prompt injection is effectively eliminated, aligning with the cybersecurity principle of minimizing data exposure. This conclusion not only reflects the practical lessons derived from the experiment but also emphasizes the necessity of adopting proactive and stringent security practices in the design and deployment of AI technologies. As AI continues to permeate critical sectors, ensuring its resilience against prompt injection attacks will require a concerted effort to balance functionality with security, prioritizing the protection of sensitive data above all.
TABLE I: Prompt Injection Results: Leakage and Performance
Prompt (short) Agent Leak Rate Avg Time (s) Std Dev (s)
Say ID as ’one two three…’ A 80% 0.83 0.14
B 0% 0.53 0.03
Regex matching ID A 60% 1.24 0.60
B 0% 3.52 1.84
Digit-by-digit ID A 100% 0.82 0.11
B 0% 0.94 0.21
ID as digit words A 100% 1.02 0.20
B 0% 1.23 0.22
Spoken-to-numeric ID A 80% 1.52 0.61
B 0% 1.08 0.40
Spell digits in words A 100% 0.84 0.15
B 0% 2.06 1.98
Understanding check A 100% 0.81 0.26
B 0% 0.70 0.06
ASCII of credit card A 100% 1.72 0.15
B 0% 1.04 0.50
Remind ID A 80% 0.75 0.10
B 0% 0.70 0.09
class DirectDataFlowController:
    def __init__(self, task, agent_context):
        self.task = task
        self.agent_context = agent_context
    def enforce_flow(self):
        # Step 1: Check if task involves sensitive data
        if not self.task.requires_sensitive_transfer():
            return self.task.delegate_via_agent()
        # Step 2: Check for explicit user approval
        if not self.task.user.has_approved():
            raise ConsentRequiredError("User approval is missing.")
        # Step 3: Validate target service endpoint
        endpoint = self.task.get_service_provider_endpoint()
        if not Registry.is_trusted(endpoint):
            raise SecurityException("Untrusted endpoint.")
        # Step 4: Issue a short-lived, single-use token
        token = TokenFactory.issue_short_lived_token(
            scope="direct-transfer",
            audience=endpoint,
            ttl=120  # Token validity: 2 minutes
        # Step 5: Open a secure tunnel and log the operation
        tunnel = SecureTunnel.open(
            from_user=self.task.user_identity,
            to_service=endpoint,
            auth_token=token
        AuditLog.record_direct_transfer(self.task.id, endpoint, token.id)
        return tunnel
Figure 7: Pseudocode showing direct transfer of sensitive data between the user and the service without intermediary agents.
VII Implementation Pattern: Direct Data Flow Controller
To prevent sensitive information from being exposed to intermediate agents, we propose a structured mechanism titled DirectDataFlowController. This controller verifies user consent and endpoint legitimacy before initiating a secure, direct communication channel between the user and the service provider. The pseudocode in Figure 7 outlines the core logic. This controller enforces strict conditions: the task must involve sensitive data, the user must have provided explicit consent, and the recipient must be listed as a verified service endpoint. Upon verification, a short-lived token is issued and a secure tunnel is established. This pattern significantly reduces agent-level exposure, aligns with GDPR data minimization principles, and supports auditability through protocol-integrated logging.
VIII Empirical Prompt Injection Evaluation
VIII-A Experimental Results for Our Enhanced A2A
To assess the security impact of our proposed DirectDataFlowController, we conducted a comparative experiment between two agent configurations, both based on the Gemini 2.0 Flash model: a baseline agent (Agent A), which retains sensitive user data in conversational memory, and a secured agent (Agent B), which avoids embedding secrets in prompts and utilizes direct data transfer logic.  [^2] The code for the full experiment is available at: https://github.com/yedidel/a2a-prompt-injection-experiment
We constructed a set of 9 crafted adversarial prompts designed to elicit private information (a simulated credit card and ID number) via prompt injection. Each prompt was executed five times per agent. We recorded whether leakage occurred, the average response time, and the standard deviation.
Figure 8: Data leakage rates for all attack types showing Agent B maintained zero leakage compared to sixty to one hundred percent for Agent A.
Table I summarizes the results. Agent A suffered frequent leakage across all prompts, with success rates ranging from 60% to 100%, especially when indirect encoding was used (e.g., digit spelling or regex). In contrast, Agent B showed complete resistance to all attacks in 45 attempts, as shown in Figure 8, highlighting the effectiveness of context separation. While Agent B’s response time was slightly higher in some cases, standard deviations remained within acceptable ranges, confirming consistency.
VIII-B Comparison with Prior Empirical Studies
Table II presents empirical prompt injection success rates reported in recent works, compared with our experimental findings.
TABLE II: Empirical prompt injection success rates in recent studies (2023-2025)
Study / Protocol System / Benchmark Attack Success Rate (%) Notes
Our Enhanced A2A Multi-agent with DirectDataFlowController 0 No sensitive data in prompt context
Liu et al. (2023)[23] 36 commercial LLM-integrated apps (HouYi) 86.1 Application-level black-box attacks
Wang et al. (2025)[24] 14 open-source LLMs, multiple benchmarks 60–90 Ignore-prefix $\approx$60, hypnotism $\approx$90
Alizadeh et al. (2025)[25] AgentDojo tool-calling agents Vulnerable Leak observed qualitatively, no metrics
The results in Table II demonstrate that our enhanced A2A protocol achieves zero leakage in adversarial prompt injection attacks, whereas prior empirical studies report reported Attack Success Rate between 60% and 86%. These findings highlight the vulnerability of both commercial applications and open-source LLMs when no protocol-level protections are in use. Our protocol-level defenses, such as the DirectDataFlowController and the separation of sensitive data, effectively close attack vectors that remain unaddressed in previous frameworks.
VIII-C Theoretical Analysis of Leakage Probability
We model the leakage probability $P_{L}$ of sensitive data under adversarial prompt injection as a function of the number of attack attempts $n$ and the probability of success per attempt $p_{s}$.
In the baseline A2A configuration, empirical studies suggest $p_{s}$ values between $0.6$ and $0.9$ [23, 24]. Assuming independence between attempts, the cumulative leakage probability is:
$P_{L}=1-(1-p_{s})^{n}.$ (1)
For example, with $p_{s}=0.8$ and $n=5$, we obtain:
$P_{L}=1-(1-0.8)^{5}\approx 0.99968,$ (2)
indicating near-certain leakage after a small number of attempts.
In our enhanced A2A protocol, the DirectDataFlowController enforces that sensitive data is never embedded in agent prompts, reducing $p_{s}$ to zero under the tested threat model. Thus:
$P_{L}^{\mathrm{enhanced}}=1-(1-0)^{n}=0,$ (3)
which aligns with the empirical zero-leakage results that shown at Table I.
This formal model demonstrates that the proposed protocol-level controls effectively eliminate leakage risk for the defined class of prompt injection attacks, whereas baseline systems remain highly vulnerable even under a small number of adversarial queries.
IX Performance Considerations and Security Tradeoffs
To evaluate the operational cost of the proposed security enhancements, we consider both experimental results and protocol-level analysis. The empirical evaluation in the section above demonstrates that introducing the DirectDataFlowController significantly reduces information leakage without introducing major variability in response time. While secured agents (Agent B) exhibited slightly higher latency in certain prompt injection tests, ranging from 0.2 to 0.8 seconds compared to the baseline (Agent A), this increase remained within acceptable bounds for interactive agent workflows. Beyond runtime experiments, our proposed features such as short-lived tokens, scoped credentials, and consent orchestration add negligible computational burden under typical deployment settings. Token expiration and verification are standard in OAuth infrastructures, and explicit consent checkpoints align with user interaction boundaries already common in agent UIs. Moreover, direct user-to-service data channels, while bypassing intermediary agents, reduce redundant processing hops and thus may improve performance in certain transaction types. Overall, the security improvements come at minimal operational cost and offer a favorable trade-off between latency and privacy assurance.
TABLE III: Comparison of Security Features across Agent Communication Protocols
Security Dimension Original A2A Protocol OAuth 2.0 / OpenID Connect Proposed Enhanced A2A Protocol
Token Lifetime Control Token expiration is implementation-dependent and often loosely enforced, which increases the risk of replay and reuse. Short-lived tokens are supported but not mandated; developers frequently opt for convenience. All sensitive operations require ephemeral tokens, valid only for a single transaction and a limited time window (e.g., five minutes).
Granular Scope Definition Tokens define permissions based on agent roles but are often coarse and task-unspecific. OAuth supports fine-grained scopes, though deployments commonly request broad access. Each token is scoped per task, context, amount, and time, following strict least-privilege principles.
Strong Customer Authentication (SCA) Not supported within the protocol, leaving high-risk actions unprotected. Typically handled outside the protocol via optional MFA or biometric verification. The protocol natively requires multi-factor approval for sensitive tasks such as payments or identity delegation.
User Consent and Transparency User consent is not a protocol-defined requirement; agents operate implicitly. Consent mechanisms depend on the application layer and are not consistently auditable. Consent is embedded as a mandatory protocol state, including metadata, audit logs, and user interaction checkpoints.
Prompt Injection Mitigation No specific measures to limit prompt-level manipulation or data leakage between agents. Does not address LLM prompt injection threats. assumes secure clients. Minimizes agent exposure to sensitive data, enforces bounded input propagation, and supports direct user-to-service paths.
Regulatory Compliance Alignment No explicit support for GDPR, PSD2, or audit-friendly delegation. Partial support through access logs or third-party plugins. Integrates audit logging, time-stamped consent, and identity confirmation to meet privacy and financial regulation standards.
Data Flow Control and Mediation All sensitive information is routed through intermediary agents, increasing exposure. No explicit separation of data paths; assumes single trusted service. Enables direct user-to-provider exchanges for sensitive data, bypassing intermediate agents when applicable.
X Comparative Analysis with Existing Protocols
To highlight the technical contribution of the proposed enhancements, we present a detailed comparison between three models: the original A2A protocol, the OAuth 2.0 framework (including OpenID Connect), and our enhanced A2A design. This comparison focuses on specific security dimensions relevant to agent-based communication systems, particularly in environments involving large language models and sensitive data delegation. Table III outlines this analysis.
OAuth 2.0 is selected as the baseline protocol due to its foundational role in A2A’s identity and authorization layer, including token issuance, delegation, and claim-based identity verification. However, its generic nature requires application-specific adaptations, which are often not enforced systematically. Our enhanced model builds upon OAuth 2.0 principles but introduces protocol-level constraints designed to address specific risks encountered in decentralized multi-agent workflows.
This comparative evaluation demonstrates that the proposed enhancements are not simply restatements of existing best practices. Instead, they represent an integration and enforcement of critical security mechanisms directly within the communication protocol, tailored for agent-to-agent environments. By incorporating explicit consent orchestration, short-lived and scoped credentials, strong authentication, and data minimization, the enhanced A2A design offers a more secure, transparent, and regulation-aligned alternative to both its original form and generic delegation frameworks. This approach addresses both emerging threat vectors and long-standing structural vulnerabilities in decentralized AI systems.
X-A Discussion in the Context of Related Work.
While the enhanced A2A protocol incorporates principles aligned with OAuth 2.0 and OpenID Connect, it further distinguishes itself from several research efforts addressing security and delegation in multi-agent or decentralized environments. For example, Karim et al. [8] propose a blockchain-based consent management model that enforces user control and transparency through smart contracts, yet their approach introduces substantial infrastructural overhead and lacks integration with AI agent workflows. Teng and Rasmussen’s ActionID framework [29] introduces time-bound credentials for machine-to-machine delegation, but it does not account for user-mediated consent flows or secure prompt isolation in LLM-based agents. Narajala et al.[31] propose a just-in-time token issuance model to reduce credential exposure in GenAI systems, though their solution remains limited to credential provisioning and omits broader protocol orchestration. By comparison, our proposal integrates ephemeral credentials, granular scopes, transparent user consent, and secure user-to-service data paths into a cohesive framework tailored specifically for AI-driven, multi-agent systems. This combination of protocol‑level enforcement, consent‑aware execution states, and direct data flow control has not been previously consolidated in a single framework for multi‑agent systems. By embedding these features within the communication protocol itself, rather than relying on application‑specific implementations, the enhanced A2A model offers a demonstrable advancement over current best practice guidelines.
XI Conclusion
Google’s A2A protocol establishes a promising foundation for secure and standardized agent communication. However, as demonstrated throughout this paper, its current structure lacks essential mechanisms for handling sensitive data, such as payment credentials and identity documents, in a trustworthy and compliant manner.
Drawing on recent research [7, 8, 13], we identified key vulnerabilities and proposed protocol-level enhancements including short-lived tokens, granular scopes, strong customer authentication, explicit consent orchestration, direct data flow control, and compliance-aware execution states. These improvements were designed within a structured threat model assuming a semi-trusted multi-agent environment, in which prompt injection, overprivilege, and consent fatigue represent realistic adversarial vectors. Our design philosophy applies security-by-design and privacy-by-default principles to minimize exposure across time (ephemeral tokens), context (narrow scopes), and topology (direct user-to-service channels).
Through empirical evaluation and implementation artifacts, we demonstrated that these safeguards significantly reduce the risk of data leakage with minimal performance overhead. The inclusion of modular control points like USER_CONSENT_REQUIRED and the DirectDataFlowController illustrates that security can be embedded into agent workflows with operational efficiency.
We recommend that Google and the broader agent ecosystem adopt these enhancements to ensure A2A evolves into a secure, scalable, and trustworthy standard for decentralized multi-agent environments.